What is the momentum of a garbage truck that is 1.40 × 10 4 1.40 × 10 4 kg and is moving at 30 m/s? p p = units At what speed would an 6 kg trash can have the same momentum as the truck? p = units = v = units =
Answer #1

Mass of Garbage truck,M=1.4104 kg

speed of truck,v=30 m/s

Momentum of truck,P=Mv=1.410430=4.2105 kg-m/s

mass of trash,m=6kg

speed of trash=P/m=(4.2105)/6=7.0104 m/s